OpenStudy (anonymous):

Let ƒ(x)=1x and g(x)=x2 − 6x. What two numbers are not in the domain of ƒ º g? Separate your answers with a comma.

OpenStudy (anonymous):

1/x not 1x typo sorry

OpenStudy (imer):

Do you mean? \[f(x)=\frac{ 1 }{ x }\]\[g(x)=x^2-6x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please help!!

OpenStudy (anonymous):

How do I figure it out? It's a timed question and I only have 2 min left

OpenStudy (imer):

\[f(g(x))=f º g(x)\], substitute g(x) in the place of "x" for f(x). Please do it for me.

OpenStudy (anonymous):

How? I'm so confused!

OpenStudy (imer):

\[f(x)=\frac{ 1 }{ x }\]where g(x) is the input. \[f(g(x))=\frac{ 1 }{ (x^2-6x) }\]

OpenStudy (anonymous):

so how do I find the answer?

OpenStudy (imer):

Now you can simplify the denominator, can you do it for me please?

OpenStudy (anonymous):

9?

OpenStudy (imer):

how 9? The denominator is \[x^2-6x\]and you see both terms have "x" common, so you re-write it as\[x(x-6)\]. We know, we can't have division by zero, for which values of "x", the denominator becomes zero?

OpenStudy (anonymous):

0.. I dont understand at all. I'm panicking because I'm running out of time and I need to get this answer correct in order to pass. Can u give me the answer and then explain to me?

OpenStudy (imer):

x=0 or x=6 is the answer.

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

and thanks so much. it was correct

OpenStudy (imer):

\[A*B=0\]is either of "A" or "B" is zero, you will get zero. \[0*B=0\]or\[A*0=0\]Now replace "A" with "x" and "B" with "x-6" and try to make each term zero individually.

OpenStudy (anonymous):

Oh! Thanks so much!!